A petrol tanker has mass 2800kg when empty and holds 30m3 of petrol when full . The density of petrol is 740kg/m3

a. Calculate the total mass of the tanker when full of petrol

b. The full tanker is travelling along a road when the driver observes an obstruction and applies the brakes. A resultant force of 30000N acts on the tanker. Calculate the deceleration of the tanker



1
Expert's answer
2021-10-21T16:33:15-0400

a)


"M=2800+(740)(30)=25000\ kg"


b)


"a=\frac{30000}{25000}=1.2\frac{m}{s^2}"